OpenStudy (anonymous):

truly need help. 2x-5y=-3 -x+2y=3

OpenStudy (anonymous):

i need to find x and y

OpenStudy (anonymous):

okay so multiply 2nd equation by 2

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

when you miltiply equation by 2 you get -2x+4y=6 and first equation is 2x-5y=-3 now add these equations you get -y=3 so y = 3

OpenStudy (anonymous):

so as x put value of y in first equation 2x-5y=-3 2x-5(-3) = -3 2x+15=-3 2x=-18 so x=-9 so y=-3 and x=-9

OpenStudy (anonymous):

in 2nd last its typo -y=3 so y=-3* ^^

OpenStudy (anonymous):

greaat :) and how can I find the value of x?

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

ok thank you sooo muchhh!!